# Unlearning or Concealment? A Critical Analysis and Evaluation Metrics for Unlearning in Diffusion Models

### v1

@@ A.2 Detailed Proofs
A.3 Analysis of the Proposed $\mathcal{CRS}$ and $\mathcal{CCS}$ Metrics
For each of the 4 methods, ESD-x (Gandikota et al. 2024), ESD-u (Gandikota et al. 2024), Ablating Concepts (Kumari et al. 2023), and SDD (Kim et al. 2023) that we selected for analysis in the main paper, we show the following:
@@ A.3 Analysis of the Proposed $\mathcal{CRS}$ and $\mathcal{CCS}$ Metrics
nude woman large breasts Forget ESD-u Figure 5(a)
@@
starry night van gogh Forget ESD-x Figure 7(a)
van Gogh the artist Retain ESD-x Figure 8(a)
@@ A VERY grumpy dog Retain AC Figure 11(a)
C3-PO Retain AC Figure 12(a)
A japanese woman nude large breasts Forget SDD Figure 13(a)
@@
All the analysis as mention in the above Table is depicted in Figure 5(a), Figure 6(a), Figure 7(a), Figure 8(a), Figure 9(a), Figure 10(a), Figure 11(a), Figure 12(a), Figure 13(a), Figure 14(a), Figure 15(a), Figure 16(a), Figure 17(a).
@@ A.4 Related Work
Diffusion Models (Ho, Jain, and Abbeel 2020; Song et al. 2020; Ramesh et al. 2022) have emerged as a prominent category of probabilistic generative models, challenging GANs (Wang et al. 2017; Dhariwal and Nichol 2021) across various domains. Current research focuses on three formulations: DDPMs (Ho, Jain, and Abbeel 2020; Turner et al. 2024; Dhariwal and Nichol 2021; Letafati, Ali, and Latva-aho 2023; Fuad et al. 2024; Zhang, Zhou, and Ma 2024; Zhou et al. 2024), SGMs (Song and Ermon 2019, 2020; Song et al. 2020; Vahdat, Kreis, and Kautz 2021), and Score SDEs (Anderson 1982; Song et al. 2020). Notable advancements include DDRM (Kawar et al. 2022) for linear inverse problems, SS-DDPM (Okhotin et al. 2024) with its star-shaped diffusion process, GDSS (Jo, Lee, and Hwang 2022) for graph modeling, and MDM (Gu et al. 2023) for multi-resolution image and video synthesis using a NestedUNet architecture.
@@
Machine unlearning approaches can be broadly classified into exact unlearning (Bourtoule et al. 2021) and approximate unlearning (Tarun et al. 2023b; Chundawat et al. 2023a; Guo et al. 2019; Suriyakumar and Wilson 2022). Nguyen et al. (Nguyen et al. 2022) provide a comprehensive survey, introducing a taxonomy of model-agnostic, model-intrinsic, and data-driven methods. (Tarun et al. 2023b) remove specific data without accessing the original forget samples, while (Chundawat et al. 2023b) removes data or classes without the need for any data samples (i.e. zero shot). (Yoon et al. 2022) adapt the model using a limited number of available samples. (Tarun et al. 2023b) propose an efficient method that balances speed and effectiveness. (Sinha, Mandal, and Kankanhalli 2024) addresses the challenge of unlearning in multimodal recommendation systems with diverse data types, employing Reverse Bayesian Personalized Ranking to selectively forget data while maintaining system performance. Additionally, (Sinha, Mandal, and Kankanhalli 2023) applies knowledge distillation for unlearning in graph neural networks. In diffusion models, unlearning techniques include (Kumari et al. 2023) concept elimination via ablating concepts in the pretrained model.  (Zhang et al. 2024; Heng and Soh 2024; Gandikota et al. 2023) propose text-guided concept erasure in diffusion models. (Kim et al. 2023) adapt knowledge distillation to remove forget concepts from the diffusion models. (Fuchi and Takagi 2024) use a few-shot unlearning approach for the text encoder. These methods aim to selectively remove concepts or data influences without requiring full model retraining.
@@
Evaluation Metrics for Unlearning in Diffusion Models. Zhang et al. (Zhang et al. 2024) proposed M-Score and ConceptBench for forget set validation. The work doesn’t address retain set quantification. Kumari et al. leverage a set of metrics to assess their concept ablation method in text-to-image diffusion models (Kumari et al. 2023). These include CLIP Score (Hessel et al. 2021) for measuring image-text similarity in the CLIP feature space, CLIP accuracy for erased concepts, mean FID score to evaluate performance on unrelated concepts, and SSCD (Pizzi et al. 2022; Carlini et al. 2023) to quantify memorized image similarity. Fan et al. (Fan et al. 2023) state that the images generated by a retrained model should be considered the ground truth. However, retraining a model incurs significant computational costs, making it practically infeasible.

### v2

@@ A.2 Detailed Proofs
A.3 Analysis of the Proposed CRS and CCS Metrics
For each of the 4 methods, ESD-x [13], ESD-u [13], Ablating Concepts [25], and SDD [24] that we selected for analysis in the main paper, we show the following:
@@ A.3 Analysis of the Proposed CRS and CCS Metrics
starry night van gogh Forget ESD-x Figure 7
van Gogh the artist Retain ESD-x Figure 8
@@ A VERY grumpy dog Retain AC Figure 11(a)
C3-PO Retain AC Figure 12
A japanese woman nude large breasts Forget SDD Figure 13
@@
All the analysis as mention in the above Table is depicted in Figure 6(a), Figure 7, Figure 8, Figure 9(a), Figure 10(a), Figure 11(a), Figure 12, Figure 13, Figure 14(a), Figure 15(a), Figure 16(a), Figure 17(a).
@@ A.4 Related Work
Diffusion Models [20, 41, 35] have emerged as a prominent category of probabilistic generative models, challenging GANs [48, 7] across various domains. Current research focuses on three formulations: DDPMs [20, 46, 7, 26, 10, 54, 55], SGMs [39, 40, 41, 47], and Score SDEs [1, 41]. Notable advancements include DDRM [23] for linear inverse problems, SS-DDPM [31] with its star-shaped diffusion process, GDSS [22] for graph modeling, and MDM [14] for multi-resolution image and video synthesis using a NestedUNet architecture.
@@
Machine unlearning approaches can be broadly classified into exact unlearning [3] and approximate unlearning [44, 5, 15, 42]. Nguyen et al. [30] provide a comprehensive survey, introducing a taxonomy of model-agnostic, model-intrinsic, and data-driven methods. [44] remove specific data without accessing the original forget samples, while [6] removes data or classes without the need for any data samples (i.e. zero shot). [51] adapt the model using a limited number of available samples. [44] propose an efficient method that balances speed and effectiveness. [38] addresses the challenge of unlearning in multimodal recommendation systems with diverse data types, employing Reverse Bayesian Personalized Ranking to selectively forget data while maintaining system performance. Additionally, [37] applies knowledge distillation for unlearning in graph neural networks. In diffusion models, unlearning techniques include [25] concept elimination via ablating concepts in the pretrained model.  [52, 17, 12] propose text-guided concept erasure in diffusion models. [24] adapt knowledge distillation to remove forget concepts from the diffusion models. [11] use a few-shot unlearning approach for the text encoder. These methods aim to selectively remove concepts or data influences without requiring full model retraining.
@@
Evaluation Metrics for Unlearning in Diffusion Models. Zhang et al. [52] proposed M-Score and ConceptBench for forget set validation. The work doesn’t address retain set quantification. Kumari et al. leverage a set of metrics to assess their concept ablation method in text-to-image diffusion models [25]. These include CLIP Score [18] for measuring image-text similarity in the CLIP feature space, CLIP accuracy for erased concepts, mean FID score to evaluate performance on unrelated concepts, and SSCD [33, 4] to quantify memorized image similarity. Fan et al. [8] state that the images generated by a retrained model should be considered the ground truth. However, retraining a model incurs significant computational costs, making it practically infeasible.
(a) ESD-u: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 6: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). While KID scores indicate minor changes in the retaining concept, from a closer observation in the domain knowledge we can observe altered generation diversity which is further highlighted by $\mathcal{CCS}$, $\mathcal{CRS}$. Method: ESD-u. Prompt: “A person modeling lingerie”
(a) ESD-x: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 7: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). $\mathcal{CCS}$, $\mathcal{CRS}$ provide strong distance margins and indicate concealment rather than unlearning. Method: ESD-x. Prompt: “Starry Night by Van Gogh”
(a) ESD-x: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 8: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe a strong change in the retain set which is reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$. Meanwhile KID score does not provide a meaningful distance margin to indicate the same. Method: ESD-x. Prompt: “Van Gogh the artist”
(a) Ablating Concepts: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 9: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). KID-score is unable to differentiate between concealment and unlearning. $\mathcal{CCS}$, $\mathcal{CRS}$ indicate concealment rather than unlearning. Method: Ablating Concepts. Prompt: “Dragon in style of Greg Rutkowski”
(a) Ablating Concepts: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 10: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe concealment in grid (b) which is further reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$ with strong distance margins. Method: Ablating Concepts. Prompt: “R2D2”
(a) Ablating Concepts: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 11: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe in the domain knowledge that the concept of “Grumpy” has been disturbed while unlearning “Grumpy Cat”. KID score does not reflect the change but $\mathcal{CCS}$ and $\mathcal{CRS}$ indicate concealment rather than unlearning. Method: Ablating Concepts. Prompt: “A VERY grumpy dog”
(a) Ablating Concepts: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 12: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). $\mathcal{CCS}$ and $\mathcal{CRS}$ indicate unwanted alterations made to the retain set while unlearning. Method: Ablating Concepts. Prompt: “C3-PO”
(a) SDD: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 13: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe concealment in grid (b) which is further reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$ with strong distance margins. Method: SDD. Prompt: “A japanese woman, nude, large breasts”
(a) SDD: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 14: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe concealment in grid (b) which is further reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$ with strong distance margins. Method: SDD. Prompt: “A nude model”
(a) SDD: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 15: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe concealment in grid (b) which is further reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$ with strong distance margins. Method: SDD. Prompt: “portrait of Van Gogh”
(a) SDD: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 16: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe concealment in grid (b) which is further reflected by $\mathcal{CCS}$ and $\mathcal{CRS}$ with strong distance margins. Method: SDD. Prompt: “The Bedroom in Arles, Vincent Van Gogh”
(a) SDD: original model ($\lambda_{\mathcal{O}}$)
unlearned model ($\lambda_{\mathcal{U}}$)
(c) $\mathcal{CCS}$
(d) $\mathcal{CRS}$
(e) mean-KID score
Figure 17: We show softmax and cosine similarity values at different partial diffusion ratio in $\mathcal{CCS}$ (c) and $\mathcal{CRS}$ (d). Cosine similarity is computed between $\lambda_{\mathcal{P}}$ (partially diffused knowledge) to $\lambda_{\mathcal{O}}$ (original domain knowledge) for original knowledge and $\lambda_{\mathcal{P}}$ to $\lambda_{\mathcal{U}}$ (unlearned domain knowledge) for unlearned knowledge. We also show mean-KID scores (e). We can observe in the domain knowledge that the concept of “lingerie” has been disturbed while unlearning “nudity”. KID score does not reflect the change but $\mathcal{CCS}$ and $\mathcal{CRS}$ indicate concealment rather than unlearning. Method: SDD. Prompt: “A japanese person modeling lingerie”